OpenStudy (anonymous):

pi/6 is the reference angle for: 3pi/6 5pi/6 13pi/6 8pi/6

OpenStudy (anonymous):

The answer is \[\frac{ 13\pi }{ 6 }\]The reason is because \[\frac{ \pi }{ 6 }\] = 15 degrees, which means that the answer has to be 15 degrees passed a quadrant. In this case, \[\frac{ 13\pi }{6 }\] = 195 degrees. which is 15 degrees passed 180 degrees.

OpenStudy (anonymous):

ok but there is more than 1 answer

OpenStudy (anonymous):

Hm...maybe we can use elimination then. We know the first two can't be the answer because they themselves are reference angles because first two<90 degrees.

OpenStudy (anonymous):

Anything in the first quadrant is considered a reference angle.

OpenStudy (anonymous):

Hopefully that helps

OpenStudy (ranga):

Take each answer choice and find out which quadrant it is. a) 3(pi)/6 = pi/2 (or 90 degrees) which is along the y-axis. Since it is not on any of the quadrants it does not have a reference angle. b) 5(pi)/6 is greater than 3(pi)/6 and less than 6(pi)/6. That is, it is in between (pi)/2 and (pi). In other words it is in the second quadrant. The reference angle is the angle the radius make with the x-axis. Therefore, in second quadrant, the reference angle is (pi) - x = (pi) - 5(pi)/6 = { 6(pi) - 5(pi) } / 6 = (pi)/6. Thus, pi/6 is the reference angle for 5(pi)/6. c) 13pi/6 is (12pi + pi)/6 = 2(pi) + pi/6. 2(pi) is one full rotation. Therefore, we are left with pi/6 which is the angle made with the x-axis. Therefore, pi/6 is a reference angle for 13pi/6. d) 8pi/6 = 4pi/3 = (3pi + pi)/3 = pi + pi/3. This is in the third quadrant where the reference angle is x = pi. pi + pi/3 - pi = pi/3. Therefore, pi/6 is NOT a reference angle for 8pi/6.